40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 176     Entry time: Thu Nov 10 17:04:05 2016
Author: Gabriele 
Type: General 
Category: Noise hunting 
Subject: Temporary second QPD 

This morning I installed temporarily a second QPD to monitor the input beam. The goal was to understand where the vibrations at frequencies below 2kHz couple from. As shown in the photo, the second QPD was close to the first one.

The signals in the two QPDs were quite different, and the coherence between them wasn't great. So I concluded that the main coupling path is not through input beam of QPD vibration, but more likely real motion of the disk.

I removed the additional QPD and restored the setup to its nominal configuration. The readout infrastructure is still in the model.

ELOG V3.1.3-